In a large population, 58% of the people have been vaccinated. If 5 people are randomly selected, what is the probability that at least one of them has been vaccinated?


1
Expert's answer
2022-04-16T04:14:07-0400

This is a binomial distribution with n=5, p=0.58.

"P(X\ge 1)=1-P(X=0)=1-(1-0.58)^{10}=0.9998."